OpenStudy (anonymous):

simplify 3aqrt(50)-2sqrt(72)+sqrt(24)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

3~(50)-2~(72)+~(24) (3*5~(2))-2~(72)+~(24) (15)~(2)-2~(72)+~(24) 15~(2)-2~(72)+~(24) 15~(2)+(-2*6~(2))+~(24) 15~(2)+(-12)~(2)+~(24) 15~(2)-12~(2)+~(24) 15~(2)-12~(2)+2~(6) 3~(2)+2~(6)

OpenStudy (anonymous):

Convert each part into surd form. 3sqrt(50) =3*(sqrt25*2) = 3* 5sqrt(2) = 15 sqrt(2) -2sqrt(72) = -2*sqrt(8*9) = -2*3sqrt(8) = -2*3*2 sqrt(2) = -12 sqrt(2) sqrt(24) = sqrt(6*4) = 2*sqrt(6) Now gather terms15 sqrt (2) -12 sqrt(2) = 3 sqrt(2) +2 sqrt(6) Answer = 3 sqrt(2) +2 sqrt(6)